# spacing=nonfrench

### v1

@@ Abstract
The exploitation of certification tools by end users represents a fundamental aspect of the development of quantum technologies as the hardware scales up beyond the regime of classical simulatability. Certifying quantum networks becomes even more crucial when the privacy of their users is exposed to malicious quantum nodes or servers as in the case of multi-client distributed blind quantum computing, where several clients delegate a joint private computation to remote quantum servers, e.g. federated quantum machine learning. In such protocols, security must be provided not only by keeping data hidden but also by verifying that the server is correctly performing the requested computation while minimizing the hardware assumptions on the employed devices. Notably, standard verification techniques fail in scenarios where the clients receive quantum states from untrusted sources such as, for example, in a recently demonstrated linear quantum network performing multi-client blind quantum computation. However, recent theoretical results provide techniques to verify blind quantum computations even in the case of untrusted state preparation. Equipped with such theoretical tools, in this work, we provide the first experimental implementation of a two-client verifiable blind quantum computing protocol in a distributed architecture. The obtained results represent novel perspectives for the verification of multi-tenant distributed quantum computation in large-scale networks.
@@
Multi-client blind quantum computing. We suppose that Alice and Bob own only single-qubit transformation devices, and want to delegate the joint computation shown in Fig. 1a to a remote quantum server while keeping their private data hidden. In our protocol, we suppose that the clients desire to perform the classical function described by the two-qubit measurement pattern $\{\phi_{1},\phi_{2}\}=\{\frac{\pi}{2},\frac{\pi}{2}\}$ where, without loss of generality, $\phi_{1}$ is chosen by Alice, while $\phi_{2}$ is chosen by Bob. Analogously, the two clients respectively insert two classical input bits, $x_{1}$ and $x_{2}$.
To mask the computation data, Alice and Bob carry out the multi-client BQC protocol introduced in [49] and depicted in Fig. 1b. However, some modifications must be applied to this protocol for the client to carry out the verification of the computation. Let us first briefly revise the non-verifiable BQC protocol introduced in [49]. Apart from the clients, the other parties involved in the protocol are (i) an untrusted source of qubits, that we will call server S1, (ii) a trusted third party (TTP) that orchestrates and secures classical communications between the clients and the server, and (iii) a remote quantum server that performs the computation, indicated as S2. In the ideal scenario where the source is behaving honestly, it distributes maximally entangled bipartite states along two Qlines, of the form:
@@
The clients sequentially apply the following one-time pad transformation to the two qubits: $R_{z}(\theta_{1}^{j})\otimes R_{z}(\theta_{2}^{j})$, where the angles $\theta_{i}^{j}$ are randomly chosen from the set $\mathcal{A}=\{0,\pi/4,...,7\pi/4\}$. We use $i$ as the qubit index and $j$ as the client index, as in Fig. 1b. Moreover, each client chooses two random bits $r_{i}^{j}$ to hide the outcome of the computation. The clients then communicate their secret parameters $\theta_{i}^{j},r_{i}^{j}$ to the TTP. We will use the following definitions: $\theta_{i}=\theta_{i}^{A}+\theta_{i}^{B}$ and $r_{i}^{j}=r_{i}^{A}\oplus r_{i}^{B}$ for the $i$-th qubit. The resulting quantum state is then sent to server S2. From now on, the clients and S2 only communicate classically, through the TTP. The protocol requires two rounds, one for each qubit, and, at the $i-$th round, the TTP computes the blind measurement angle $\delta_{i}$ according to the formula:
@@
where $m_{0}^{\text{true}}=0$ and $m_{1}^{\text{true}}=m_{1}\oplus r_{1}$. The outcome of the computation is finally corrected by the TTP according to $m_{i}^{\text{true}}=m_{i}\oplus r_{i}$. The protocol is repeated many times and the outcome of the computation is chosen through majority vote among all rounds.
@@
A second modification is due to the need of changing the state preparation procedure. In detail, the clients now need to randomly prepare the received qubits according to the transformation $R_{z}(\theta_{1}^{j})X^{b_{1}^{j}}\otimes R_{z}(\theta_{2}^{j})X^{b_{2}^{j}}$, as shown in Fig. 1c. The bits $b_{i}^{j}$ are randomly chosen by the clients to perform the verification protocol [57]. In computation rounds, the blind measurement angle $\delta_{i}$ at the $i-$th round is computed by the TTP according to the formula:
$\delta_{i}=\theta_{i}^{\prime}+x_{i}\pi+(-1)^{m_{i-1}^{\text{true}}}\phi_{i}+r_{i}\pi$ (3)
where $m_{0}^{\text{true}}=0$, $m_{1}^{\text{true}}=m_{1}\oplus r_{1}$, and $\theta^{\prime}_{i}=(-1)^{b_{i}^{B}}\theta_{i}^{A}+\theta_{i}^{B}+(b_{i}^{A}\oplus b_{i}^{B})\pi$. In test rounds, instead, the target algorithm of the clients is the stabilizer measurement $Y\otimes Y$. Indeed, since the outcome of such measurements is known, in this way the clients can detect eventual malicious deviation on the server side [41, 57]. In practice, this means that the second measurement is not chosen adaptively and, therefore, $m_{2}^{\text{true}}$ is set to 00 for both qubits. At the end of the protocol, the outcomes are decrypted according to the formula: $m_{i}^{\text{true}}=m_{i}\oplus r_{i}$.
@@
Figure 2: Experimental setup. A Sagnac-based source generates entangled photon pairs in the state in Eq. (1). The pairs are sent via a fiber link to the two clients who perform their transformation through HWPs and LCs. The entangled pair is then sent to server $S_{2}$, which measures the two qubits. The second qubit measurement is chosen according to the first measurement outcome via an active feed-forward system acting on an electro-optical modulator, i.e. a Pockels cell (PC).
@@
Experimental setup. In Fig. 2 we describe the experimental apparatus employed to implement the two-client protocol. A Sagnac-based source of polarization-entangled photons [58], i.e. server S1, generates pairs of photons in the state defined in Eq. (1), where the state vector |$\ket{0}$ is encoded in the photons horizontal polarization (|$\ket{H}$), and |$\ket{1}$ in the vertical one (|$\ket{V}$). The photon pairs are generated in a spontaneous parametric down-conversion-based source, composed of a 25-mm-thick periodically-poled Potassium Titanyl Phosphate (ppKTP) crystal inside a Sagnac interferometer [58]. Photon pairs are sent to the clients who sequentially apply their random transformations through half-wave plates (HWP) rotated at $0^{\circ}$ or $45^{\circ}$, respectively depending on the bit $b_{i}^{j}$, and liquid crystals (LC). At each run, the clients’ random parameters are chosen through an ID Quantique quantum random number generator (QRNG). The QRNG is also used to choose run-by-run between test and computational rounds. The TTP is made up of a computer equipped with an input/output digital module and stores all clients’ parameters. With such information, the TTP computes the measurement angle $\delta_{1}$, and the first measurement station is set accordingly. Moreover, the TTP pre-computes the two possible values for $\delta_{2}$, namely $\delta_{2}^{\pm}=\theta^{\prime}_{2}+x_{2}\pi+r_{2}\pi\pm\phi_{2}$ for computation rounds or $\delta_{2}=\theta^{\prime}_{2}+x_{2}\pi+r_{2}\pi+\phi_{2}$ for test rounds, where $\phi_{1}=\phi_{2}=\pi/2$. The two photons are sent to server S2 that performs measurements of the form $M(\delta)=\cos{(\delta)}\sigma_{x}+\sin{(\delta)}\sigma_{y}$. The first measurement station is made up of a quarter-wave plate (QWP), a half-wave plate (HWP), and a polarizing beam splitter (PBS). The two single-photon avalanche photodiodes of the first measurement station are connected to a fast electronic circuit that, in computation rounds, selects $\delta_{2}^{+}$ or $\delta_{2}^{-}$, according to the corrected outcome of the first measurement, $m_{1}^{\text{true}}=m_{1}\oplus r_{1}$. In test rounds, instead, the second measurement station is fixed to $\delta_{2}^{+}$. In the second measurement station, we substitute the QWP with a Pockels cell (PC), i.e. a fast electro-optical modulator that performs the identity when no voltage is applied, while applying a phase shift between orthogonal polarizations when a voltage is applied. The second photon is delayed by using a $\approx 65$ m single-mode fiber to enable active feed-forward in the second measurement station. Finally, the second outcome is corrected according to $m_{2}^{\text{true}}=m_{2}\oplus r_{2}$.
@@
Figure 3: Experimental results. a)-b) Theoretical and experimental outcome probabilities of the test and computation rounds, respectively, for the algorithm $\left(\phi_{1},\phi_{2},x_{1},x_{2}\right)=\left(\frac{\pi}{2},\frac{\pi}{2},0,0\right)$. c)-d) Theoretical and experimental outcome probabilities of the test and computation rounds, respectively, for the algorithm $\left(\phi_{1},\phi_{2},x_{1},x_{2}\right)=\left(\frac{\pi}{2},\frac{\pi}{2},1,0\right)$. e) Robustness and security of the protocol. We find that, for both algorithms, the experimental test error fraction $\epsilon$ is such that $\nu<\epsilon<\omega$. Therefore, the protocol shows robustness and security.
@@
We also checked for protocol blindness. In detail, in Fig. 4a, we show the density matrix obtained by averaging over the $64$ density matrices obtained by combining all possible clients’ rotation angles $\theta_{1}^{A}$ and $\theta_{1}^{B}$ and performing quantum state tomography of the second qubit, while the first qubit is measured in the basis $\delta_{1}=5\pi/4$. The fidelity with the completely mixed state amounts to $F_{1q}=0.9952\pm 0.0003$. In Fig. 4b, instead, we show the density matrix obtained by averaging over the $64$ density matrices obtained by combining all possible clients’ rotation angles $\theta_{1}^{A}$ and $\theta_{2}^{B}$. The fidelity with the completely mixed state amounts to $F_{2q}=0.9901\pm 0.0002$. To quantify the classical information that the server can retrieve from the received full two-qubit quantum state, we compute the Holevo quantity $\chi$ of the two-qubit state. Such a quantity is defined as:

### v2

@@ Abstract
The exploitation of certification tools by end users represents a fundamental aspect of the development of quantum technologies as the hardware scales up beyond the regime of classical simulatability. Certifying quantum networks becomes even more crucial when the privacy of their users is exposed to malicious quantum nodes or servers as in the case of multi-client distributed blind quantum computing, where several clients delegate a joint private computation to remote quantum servers, such as federated quantum machine learning. In such protocols, security must be provided not only by keeping data hidden but also by verifying that the server is correctly performing the requested computation while minimizing the hardware assumptions on the employed devices. Notably, standard verification techniques fail in scenarios where the clients receive quantum states from untrusted sources such as, for example, in a recently demonstrated linear quantum network performing multi-client blind quantum computation. However, recent theoretical results provide techniques to verify blind quantum computations even in the case of untrusted state preparation. Equipped with such theoretical tools, in this work, we provide the first experimental implementation of a two-client verifiable blind quantum computing protocol in a distributed architecture. The obtained results represent novel perspectives for the verification of multi-tenant distributed quantum computation in large-scale networks.
@@
Multi-client blind quantum computing. We suppose that Alice and Bob own only single-qubit transformation devices, and want to delegate the joint computation shown in Fig. 1a to a remote quantum server while keeping their private data hidden. In our protocol, we suppose that the clients desire to perform the classical function described by the two-qubit measurement pattern $\{\phi_{1},\phi_{2}\}=\{\frac{\pi}{2},\frac{\pi}{2}\}$. The two clients respectively insert two classical input bits, $x_{1}$ and $x_{2}$. To mask the computation data, Alice and Bob carry out the multi-client BQC protocol introduced in [49] and depicted in Fig. 1b. However, some modifications must be applied to this protocol for the client to carry out the verification of the computation. Let us first briefly revise the non-verifiable BQC protocol introduced in [49]. Apart from the clients, the other parties involved in the protocol are (i) an untrusted source of qubits, that we will call server S1, (ii) a trusted third party (TTP) that orchestrates and secures classical communications between the clients and the server, and (iii) a remote quantum server that performs the computation, indicated as S2. In the ideal scenario where the source is behaving honestly, it distributes maximally entangled bipartite states along two Qlines, of the form:
@@
[…]
@@
where $m_{0}^{\text{true}}=0$ and $m_{1}^{\text{true}}=m_{1}\oplus r_{1}$. Both outcomes are corrected by the TTP according to $m_{i}^{\text{true}}=m_{i}\oplus r_{i}$ at the end of the protocol. The protocol is repeated many times and the outcome of the computation is chosen through majority vote among all rounds.
@@
A second modification is due to the need of changing the state preparation procedure. In detail, the clients now need to randomly prepare the received qubits according to the transformation $R_{z}(\theta_{1}^{j})X^{b_{1}^{j}}\otimes R_{z}(\theta_{2}^{j})X^{b_{2}^{j}}$, as shown in Fig. 1c. The bits $b_{i}^{j}$ are randomly chosen by the clients to perform the verification protocol [57] and secretly communicated to the TTP. In computation rounds, the blind measurement angle $\delta_{i}$ at the $i-$th round must be computed in an adaptive way by the TTP according to the formula given in Eq. (2), where the angles $\theta_{i}$ are substituted with the angles $\theta^{\prime}_{1}=(-1)^{b_{1}^{B}}\theta_{1}^{A}+\theta_{1}^{B}+(b_{2}^{A}\oplus b_{2}^{B})\pi$ while $\theta^{\prime}_{2}=(-1)^{b_{2}^{B}}\theta_{2}^{A}+\theta_{2}^{B}+(b_{1}^{A}\oplus b_{1}^{B})\pi$. In test rounds, instead, the target algorithm of the clients is the stabilizer measurement $Y\otimes Y$. Indeed, since the outcome of such measurements is known, in this way the clients can detect eventual malicious deviation on the server side [41, 57]. In practice, this means that the second measurement is fixed to $\delta_{2}=\theta_{2}^{\prime}+x_{2}\pi+\phi_{2}+r_{2}\pi$.
@@
Figure 2: Experimental setup. A Sagnac-based source generates entangled photon pairs in the state in Eq. (1). The pairs are sent via a fiber link to the two clients who perform their transformation through HWPs and LCs. The entangled pair is then sent to server $S_{2}$, which measures the two qubits. The second qubit measurement is chosen according to the first measurement outcome via an active feed-forward system acting on an electro-optical modulator, i.e. a Pockels cell.
@@
Experimental setup. The experimental setup employed is depicted in Fig. 2. Server S1 is a Sagnac-based source of polarization-entangled photon pairs [58]. The qubit state |$\ket{0}$ is encoded in the photons’ horizontal polarization (|$\ket{H}$) and |$\ket{1}$ in the vertical one (|$\ket{V}$). The TTP is made up of a computer equipped with an input/output digital module and a fast digital circuit. Further information about the feed-forward system can be found in the Supplementary Information. The clients’ secret parameters are stored in the TTP computer and used to compute the first measurement basis, $\delta_{1}$, and the two possible values of the second measurement basis, $\delta_{2}$ for computation rounds, i.e. $\delta_{2}^{\pm}=\theta^{\prime}_{2}+x_{2}\pi+r_{2}\pi\pm\phi_{2}$. In test rounds, the second measurement basis is fixed to $\delta_{2}^{+}$. Server S2 is made up of two polarization measurement stations that carry out measurements of the form $M(\delta)=\cos{(\delta)}\sigma_{x}+\sin{(\delta)}\sigma_{y}$. The first measurement station uses a quarter-wave plate (QWP), a half-wave plate (HWP), and a polarizing beam splitter (PBS). In the second measurement station, the QWP is substituted with a fast electro-optical modulator, i.e. a Pockels cell (PC), which is activated in $\approx 325$ ns with high-voltage pulses, while the second photon is delayed through a $\approx 65$ m single-mode fiber. The avalanche photodiodes (APDs) in the first measurement station are linked to the TTP digital circuit to enable adaptivity of the second measurement basis. The clients apply their random transformations through half-wave plates rotated at $0^{\circ}$ or $45^{\circ}$, depending on the bit $b_{i}^{j}$, and liquid crystals (LC). The clients’ random parameters are chosen through an ID Quantique quantum random number generator (QRNG). The QRNG is also used to choose run-by-run between test and computational rounds.
@@
Figure 3: Experimental results. a)-b) Theoretical and experimental outcome probabilities of the test and computation rounds, respectively, for the algorithm $\left(\phi_{1},\phi_{2},x_{1},x_{2}\right)=\left(\frac{\pi}{2},\frac{\pi}{2},0,0\right)$. c)-d) Theoretical and experimental outcome probabilities of the test and computation rounds, respectively, for the algorithm $\left(\phi_{1},\phi_{2},x_{1},x_{2}\right)=\left(\frac{\pi}{2},\frac{\pi}{2},1,0\right)$. e) Robustness and security of the protocol. We find that, for both algorithms, the experimental test error fraction $\epsilon$ is such that $\epsilon<\omega$. Therefore, the protocol shows robustness and security.
@@
We also checked for protocol blindness. In detail, in Fig. 4a, we show the density matrix obtained as an average over $64$ combinations, namely all possible clients’ rotation angles $\theta_{1}^{A}$ and $\theta_{1}^{B}$, and performing quantum state tomography of the second qubit, while the first qubit is measured in the basis $\delta_{1}=5\pi/4$. The fidelity with the completely mixed state amounts to $F_{1q}=0.9952\pm 0.0003$. In Fig. 4b, instead, we show the density matrix obtained by averaging over the $64$ density matrices obtained by combining all possible clients’ rotation angles $\theta_{1}^{A}$ and $\theta_{2}^{B}$. The fidelity with the completely mixed state amounts to $F_{2q}=0.9901\pm 0.0002$. To quantify the classical information that the server can retrieve from the received full two-qubit quantum state, we compute the Holevo quantity $\chi$ of the two-qubit state. Such a quantity is defined as: